# Analog Quantum Simulation of Coupled Electron-Nuclear Dynamics in Molecules

### v2

@@ Abstract
Quantum computing has the potential to reduce the computational cost required for quantum dynamics simulations. However, existing quantum algorithms for coupled electron-nuclear dynamics simulation either require fault-tolerant devices, or use the Born-Oppenheimer (BO) approximation and a truncation of the electronic basis. We present the first analog quantum simulation approach for molecular vibronic dynamics in a pre-BO framework, i.e. without the separation of electrons and nuclei, by mapping the molecular Hamiltonian to a device with coupled qubits and bosonic modes. We show that our approach has exponential savings in resource and computational costs compared to the equivalent classical algorithms. The computational cost of our approach is also exponentially lower than existing BO-based quantum algorithms. Furthermore, our approach has a much smaller resource scaling than the existing pre-BO quantum algorithms for chemical dynamics. The low cost of our approach will enable an exact treatment of electron-nuclear dynamics on near-term quantum devices.
@@ I Introduction
Light-matter interactions are the source of many phenomena in molecular systems such as vision [1, 2, 3], photosynthesis [4, 5, 6, 7], photovoltaics [8, 9, 10, 11], and photocatalysis [12, 13]. In order to apply light-matter interactions to the development of high-functional molecular devices, it is important to understand their underlying mechanisms. In principle, the molecular time-dependent Schrödinger equation can be exactly solved for systems of electrons and nuclei to simulate molecular dynamics. However, such a solution is practically impossible for molecules with more than a few atoms due to the exponential scaling of the computational cost with respect to system size.
The Born-Oppenheimer (BO) approximation neglects the coupling between electronic and nuclear degrees of freedom, known as nonadiabatic coupling (NAC) [14]. While it is useful for properties of the ground electronic state, it fails to describe light-induced chemical reactions where the molecule can reach strong NAC regions after photo-excitation of the molecular electronic state [14]. The BO approximation can be extended to the group BO approximation (GBOA), which takes into account NACs between a group of BO states considered to be relevant to the dynamics, while NACs to states outside of the group are neglected [14]. The molecular dynamics are then described as nuclear dynamics on multiple BO potential energy surfaces coupled via NACs. The majority of existing simulation methods make use of the GBOA since it extends the scope of molecular dynamics simulation while using conventional electronic structure methods [15, 16, 17]. However, truncation of the electronic basis by the GBOA requires some degree of intuition, and it necessarily neglects couplings to higher-lying BO states that alter the molecular dynamics. Furthermore, since existing simulation methods with the GBOA require an accurate calculation of BO states and their gradients (including NACs), the computational cost for accurate simulations can become intractable.
@@
In this work, we propose the first analog quantum simulation method for molecular vibronic dynamics in the pre-BO framework. Our approach maps the second quantized pre-BO representation of the molecular vibronic Hamiltonian onto an analog quantum device, and thus treats the electron-nuclear interactions exactly. Specifically, we map the nuclear vibrational motions to bosonic modes of a device and use a fermion-qubit mapping for electronic degrees of freedom. We show that our method can efficiently simulate the exact molecular vibronic dynamics within a given single-particle basis set, and therefore it is suitable for accurate near-term quantum simulations of coupled electron-nuclear dynamics. As an example, we show how a single-mode vibronic model system could be simulated on currently available trapped ion quantum computers with existing experimental techniques and noise. Finally, we show how the scaling of our approach out-performs all previous approaches in terms of hardware resources, implementation cost, and classical pre-calculation.
@@ II.1 Theory
Our simulation approach is restricted to vibronic (vibrational + electronic) internal degrees of freedom of the molecule, meaning the translational and rotational degrees of freedom of the atomic nuclear coordinates are removed. In general, this can be achieved by any unitary transformation of the Cartesian atomic coordinates that separates the 6 collective translations and rotations of the molecule, given by vectors $\mathbf{Q}_{\mathrm{trans}}$ and $\mathbf{Q}_{\mathrm{rot}}$, from the $63N_{\mathrm{at}}-$ 6 vibrational internal coordinates, $\mathbf{Q}\equiv\mathbf{Q}_{\mathrm{vib}}$. For the remainder of the manuscript, we assume that mass-weighted normal mode coordinates are used, whereby the translations and rotations are easily identified as the zero-frequency modes. The electronic coordinates are defined as positions of electrons relative to the nuclear center of mass with a fixed orientation, i.e. in the Eckart frame [45]. The Coriolis coupling (between nuclear vibrations and rotations) is excluded. While the removal of molecular rotations is an approximation [45], it is appropriate for ultrafast (fs–ps) chemical dynamics simulations, as is further justified in the Discussion.
@@
We express the molecular vibronic wavefunction using Slater determinants of the position-dependent spin orbitals for the electronic degrees of freedom and Hartree products of harmonic oscillator eigenstates of each normal mode for the vibrational degrees of freedom. In the second quantization representation, the Slater determinants and the Hartree product can be expressed as occupation number vectors (ONVs). An electronic ONV for a Slater determinant is written as $|\mathbf{n}\rangle_{\mathrm{e}}=|n_{1}\cdots n_{N_{\mathrm{o}}}\rangle_{\mathrm{e}}$, where $n_{j}$ is the occupation of the $j$-th spin orbital and $N_{\mathrm{o}}$ is the number of spin orbitals, while a vibrational ONV for a Hartree product is given as $|\mathbf{v}\rangle_{\mathrm{n}}=|v_{1}\cdots v_{N_{\mathrm{mode}}}\rangle_{\mathrm{n}}$, where $v_{\nu}$ is the occupation (Fock state) of mode $\nu$. Therefore, our wavefunction ansatz becomes
@@
$\begin{split}\hat{H}_{\mathrm{mol}}&=\sum_{pq}^{N_{\mathrm{o}}}h_{pq}(\hat{\mathbf{Q}})\hat{a}^{\dagger}_{p}\hat{a}_{q}+\dfrac{1}{2}\sum_{pqrs}^{N_{\mathrm{o}}}v_{pqrs}(\hat{\mathbf{Q}})\hat{a}^{\dagger}_{p}\hat{a}^{\dagger}_{q}\hat{a}_{s}\hat{a}_{r}\\ &+\sum_{pq}^{N_{\mathrm{o}}}\sum_{\nu}^{N_{\mathrm{mode}}}\left(id_{\nu,pq}(\hat{\mathbf{Q}})\cdot\hat{P}_{\nu}-g_{\nu,pq}(\hat{\mathbf{Q}})\right)\hat{a}^{\dagger}_{p}\hat{a}_{q}\\ &+\sum_{\nu}^{N_{\mathrm{mode}}}\dfrac{\hat{P}_{\nu}^{2}}{2}+V_{\mathrm{nn}}(\hat{\mathbf{Q}}),\end{split}$ (3)
@@
For fixed Taylor expansion orders, $\omega_{\nu}$ and $\mathbf{Q}_{0}$ only affect the accuracy of our approach by the number of harmonic eigenstates required to describe the wavefunction. In other words, a quantum simulator with lower oscillator noise can tolerate lower-accuracy parameters, including approximations derived from the electronic integral Taylor expansions. For practical cases in the near term, we expect that density functional theory will provide an ideal balance between accuracy and classical computational cost when finding $\mathbf{Q}_{0}$ and $\omega_{\nu}$, and Kohn-Sham orbitals provide a compact spin orbital basis [47].
@@
Figure 1: Illustration of the pre-BO cMQB analog quantum simulation approach for molecular vibronic dynamics using trapped-ion architectures in comparison with the BO framework. (a) The electronic wavefunction is given in terms of spin orbitals. Each diabatic spatial orbital, which depends on the nuclear positions, yields two spin orbitals with opposite spins: $|\phi_{1}\rangle$ and $|\phi_{3}\rangle$ (green), and $|\phi_{2}\rangle$ and $|\phi_{4}\rangle$ (yellow). Only the motion of a single ion (grey) is shown. (b) Molecular vibronic dynamics are simulated based on the pre-BO representation with interactions among electronic (cyan) and nuclear (grey) degrees of freedom. (c) In this work, we propose a coupled multi-qubit boson approach which maps occupation numbers $\{n_{p}\}$ of spin orbitals $\{|\phi_{p}\rangle\}$ to qubits and nuclear motions to bosonic modes coupled to the qubits, and thus can perform pre-BO dynamics quantum simulations with a complete electronic basis within a given orbital basis set. (d) Existing quantum/classical algorithms rely on BO states $\{|\Phi_{i}\rangle\}$ with potential energy surfaces $\{|E_{i}\rangle\}$ found by solving the electronic time-independent Schrödinger equation using the orbital basis at fixed nuclear positions prior to the simulation. (e) Molecular vibronic dynamics are solved in the BO framework by propagating the nuclear wavefunction (red and blue densities) on a truncated set of coupled BO states. The truncation leads to a deviation from exact results shown in (b).
@@ II.2 Analog quantum simulation
We now introduce our approach to map the Hamiltonian $\hat{H}_{\mathrm{mol}}$ onto an analog quantum device to perform a pre-BO simulation. Such a mapping must satisfy the symmetry requirements of fermions (electrons) and bosons (vibrations), given by their (anti-)commutation relations. Ideally, we would like to map both types of degrees of freedom onto a device with native fermionic [50, 51] and bosonic degrees of freedom. However, there is currently no practical quantum architecture that contains both degrees of freedom with individual control. Therefore, we will target quantum architectures that have qubit levels and bosonic modes with controllable coupling between them, such as ion traps and circuit quantum electrodynamics (cQED), using a fermion-qubit mapping [52, 53]. We will refer to such architectures as “coupled multi-qubit-boson” (cMQB) devices.
An illustration of our pre-BO analog simulation method in comparison with the BO framework is summarized in Fig. 1. We start by finding a nuclear position dependent spin orbital basis [Fig. 1(a)] and deriving the expansion coefficients in Eq. (3). We then encode the pre-BO wavefunction [Fig. 1(b)] on a cMQB device, e.g. a trapped-ion device with ion electronic states representing electronic ONVs, and ion motional modes representing the nuclear component of the wavefunction [Fig. 1(c)]. In contrast, existing classical/quantum algorithms in the BO framework require the pre-calculation of BO electronic states [Fig. 1(d)] with the spin orbital basis. The vibronic wavefunction is then propagated on a truncated BO basis [Fig. 1(e)], which can result in inaccurate dynamics simulations.
@@
The direct implementation of our approach outlined above requires the simultaneous control of all degrees of freedom in the simulation. However, our approach involves a fermion-qubit mapping on the electronic degrees of freedom, which involves a change of basis between non-commuting electronic terms. Therefore, we must use Trotterization [54] of the time-evolution operator. Trotterization approximates the time-evolution operator as a product of operators, each of which corresponds to a term $\hat{H}_{I}$ in Eq. (4). The Trotterized cMQB time-evolution operator is given by $\hat{U}(t)\approx\left(\prod^{N_{\mathrm{op}}}_{I}\hat{U}_{I}(t/N_{\mathrm{t}})\right)^{N_{\mathrm{t}}}\equiv\hat{U}_{\mathrm{cMQB}}(\Delta t)^{N_{\mathrm{t}}}$, where $N_{\mathrm{op}}$ is the number of divided operators and $N_{\mathrm{t}}$ is the number of the Trotter steps [Fig. 2(b)].
@@
where $\hat{\Lambda}^{I}$ is an $N$-qubit Clifford operation which can be composed out of Hadamard, phase, and CNOT gates, to propagate the coupling between bosonic mode(s) and a single qubit $q_{0}^{I}$ generated by a laser-ion interaction to multiple qubits for the desired Pauli strings for $\hat{H}_{I}$. In Fig. 2(b) we show an example time-evolution operator $\hat{U}_{I}(\Delta t)=\exp(-i\theta\Delta t(\hat{b}_{1}+\hat{b}_{1}^{\dagger})\hat{Y}_{2}\hat{Z}_{3}\hat{Y}_{4})$, which can be achieved by a sequence of digital and analog quantum operations. The nuclear-only terms of the Hamiltonian (the final two terms of Eq. (3) are given up to second order by a constant detuning from the bosonic mode frequency. Higher-order nuclear-only terms can be achieved with operations on only the bosonic modes, or qubit-boson operations with an ancilla qubit. This approach can be considered a digital-analog quantum simulation [55, 56] because the gates involved in generating qubit entanglement take a digital form. A fully analog form of the time-evolution could be achieved using time- and spin-dependent squeezing/displacement operators on the collective ionic motions [57]; however, for the remainder of this paper we consider the digital-analog approach.
@@
Trapped ions and cQED typically use bosonic modes to achieve entangling gates in digital quantum algorithms. The quantum gates necessary to implement $\hat{\Lambda}^{I}$ are thus readily achieved with these architectures. However, since the entangling gates require a bosonic mode, some bosonic modes must be reserved from those used for simulating nuclear vibrational degrees of freedom. We refer to these as the bus modes. The total number of bosonic modes in the simulator must therefore be at least one greater than the total number of molecular vibrational modes.
@@
As an analog simulation approach, our approach is subject to environmental noise over the course of the simulation. The noise effect on an open quantum system can be described by the Lindblad master equation [60]:
@@
where $\hat{\rho}=|\Psi\rangle\langle\Psi|$ represents the density operator of the system, $\hat{H}$ is the Hamiltonian of the system, $D[\hat{L}]\hat{\rho}=\hat{L}\hat{\rho}\hat{L}^{\dagger}-\frac{1}{2}\{\hat{L}^{\dagger}\hat{L},\hat{\rho}\}$ is the Lindblad superoperator, $\hat{L}_{i}$ is the jump operator for the noise source $i$, and $\gamma_{i}$ is the corresponding dissipation rate. The noise effect on a cMQB device translates into the molecular state during a cMQB simulation, where the native rates of the cMQB simulator are scaled to the molecular system depending on the encoding, device parameters, and simulation setup.
@@
where $\Delta t_{\mathrm{mol}}$ is the Trotter step size in the molecular scale, $t_{\mathrm{CNOT}}$ is the experimental time for a two-qubit entangling gate, and $N_{\mathrm{CNOT}}$ is the number of two-qubit entanglement gates per single Trotter step (See Supplemental Material [46]). Although a small Trotter step $\Delta t_{\mathrm{mol}}$ reduces the theoretical error of the cMQB time-evolution operator, the above equations imply that trade-off exists due to the number of digital qubit entanglement operations.

### v3

@@ Abstract
Quantum computing has the potential to reduce the computational cost required for quantum dynamics simulations. However, existing quantum algorithms for coupled electron-nuclear dynamics simulation either require fault-tolerant devices, or involve the Born-Oppenheimer (BO) approximation and pre-calculation of electronic states on classical computers. We present the first quantum simulation approach for molecular vibronic dynamics in a pre-BO framework with an analog mapping of nuclear degrees of freedom, i.e. without the separation of electrons and nuclei, by mapping the molecular Hamiltonian to a device with coupled qubits and bosonic modes. We perform a proof-of-principle emulation of our ansatz using a single-mode model system which represents vibronic dynamics of chemical systems, such as nonadiabatic charge transfer involving polarization of the medium, and propose an implementation of our approach on a trapped-ion device. We show that our approach has exponential savings in resource and computational costs compared to the equivalent classical algorithms. Furthermore, our approach has a much smaller resource and implementation scaling than the existing pre-BO quantum algorithms for chemical dynamics. The low cost of our approach will enable an exact treatment of electron-nuclear dynamics on near-term quantum devices.
@@ I Introduction
Light-matter interactions are the source of many phenomena in molecular systems such as vision [1, 2, 3], photosynthesis [4, 5, 6, 7], photovoltaics [8, 9, 10, 11], and photocatalysis [12, 13]. In order to apply light-matter interactions to the development of high-functional molecular devices, it is important to understand their underlying mechanisms. Such interactions often involve nuclear quantum effects that can dramatically alter reaction dynamics relative to classical predictions. In principle, the molecular time-dependent Schrödinger equation can be exactly solved for systems of electrons and nuclei to simulate molecular dynamics. However, such a solution is practically impossible for molecules with more than a few atoms due to the exponential scaling of the computational cost with respect to system size.
The Born-Oppenheimer (BO) approximation neglects the coupling between electronic and nuclear degrees of freedom, known as nonadiabatic coupling (NAC) [14]. While it is useful for properties of the ground electronic state, it fails to describe light-induced chemical reactions where the molecule can reach strong NAC regions after photo-excitation of the molecular electronic state [14]. The BO approximation can be extended to the group BO approximation (GBOA), which takes into account NACs between a group of BO states considered to be relevant to the dynamics, while NACs to states outside of the group are neglected [14]. The molecular dynamics are then described as nuclear dynamics on multiple BO potential energy surfaces coupled via NACs. The majority of existing simulation methods make use of the GBOA since it extends the scope of molecular dynamics simulation while using conventional electronic structure methods [15, 16, 17]. Ideally, the couplings neglected by the GBOA are minimal, but choosing the number of BO electronic states requires some intuition. Furthermore, since existing simulation methods with the GBOA require an accurate calculation of BO states and their gradients (including NACs), the computational cost for accurate simulations can become intractable.
@@
In this work, we propose the first quantum simulation method for molecular vibronic dynamics in the pre-BO framework with an analog mapping of nuclear degrees of freedom. Our approach maps the second quantized pre-BO representation of the molecular vibronic Hamiltonian onto an analog quantum device, and thus treats the electron-nuclear interactions exactly. Specifically, we map the nuclear vibrational motions to bosonic modes of a device and use a fermion-qubit mapping for electronic degrees of freedom. We show that our method can efficiently simulate the exact molecular vibronic dynamics within a given single-particle basis set, and therefore it is suitable for accurate near-term quantum simulations of coupled electron-nuclear dynamics. As an example, we show how a single-mode vibronic model system of nonadiabatic charge transfer [45, 46] could be simulated on currently available trapped ion quantum computers with existing experimental techniques and noise. Finally, we show how our approach has the potential to out-perform all previous approaches in terms of scaling of hardware resources, implementation cost, and classical pre-calculation.
@@ II.1 Theory
Our simulation approach is restricted to vibronic (vibrational + electronic) internal degrees of freedom of the molecule, meaning the translational and rotational degrees of freedom of the atomic nuclear coordinates are removed. In general, this can be achieved by any unitary transformation of the Cartesian atomic coordinates that separates the 6 collective translations and rotations of the molecule, given by vectors $\mathbf{Q}_{\mathrm{trans}}$ and $\mathbf{Q}_{\mathrm{rot}}$, from the $3N_{\mathrm{at}}-6$ vibrational internal coordinates, $\mathbf{Q}\equiv\mathbf{Q}_{\mathrm{vib}}$. For the remainder of the manuscript, we assume that mass-weighted normal mode coordinates are used, whereby the translations and rotations are easily identified as the zero-frequency modes. The electronic coordinates are defined as positions of electrons relative to the nuclear center of mass with a fixed orientation, i.e. in the Eckart frame [47]. The Coriolis coupling (between nuclear vibrations and rotations) is excluded. While the removal of molecular rotations is an approximation [47], it is appropriate for ultrafast (fs–ps) chemical dynamics simulations, as is further justified in the Discussion.
@@
We express the molecular vibronic wavefunction using Slater determinants of the position-dependent spin orbitals for the electronic degrees of freedom, and Hartree products of harmonic oscillator eigenstates of each normal mode for the vibrational degrees of freedom. In the second quantization representation, the Slater determinants and the Hartree product can be expressed as occupation number vectors (ONVs). An electronic ONV for a Slater determinant is written as $|\mathbf{n}\rangle_{\mathrm{e}}=|n_{1}\cdots n_{N_{\mathrm{o}}}\rangle_{\mathrm{e}}$, where $n_{j}$ is the occupation of the $j$-th spin orbital and $N_{\mathrm{o}}$ is the number of spin orbitals, while a vibrational ONV for a Hartree product is given as $|\mathbf{v}\rangle_{\mathrm{n}}=|v_{1}\cdots v_{N_{\mathrm{mode}}}\rangle_{\mathrm{n}}$, where $v_{\nu}$ is the occupation (Fock state) of mode $\nu$. Therefore, our wavefunction ansatz becomes
@@
[…]
@@
[…]
@@
Figure 1: Illustration of the pre-BO cMQB analog quantum simulation approach for molecular vibronic dynamics using trapped-ion architectures in comparison with the BO framework. (a) The electronic wavefunction is given in terms of spin orbitals. Each diabatic spatial orbital, which depends on the nuclear positions, yields two spin orbitals with opposite spins: $|\phi_{1}\rangle$ and $|\phi_{3}\rangle$ (green), and $|\phi_{2}\rangle$ and $|\phi_{4}\rangle$ (yellow). Only the motion of a single ion (grey) is shown. (b) Molecular vibronic dynamics are simulated based on the pre-BO representation with interactions among electronic (cyan) and nuclear (grey) degrees of freedom. (c) In this work, we propose a coupled multi-qubit boson approach which maps occupation numbers $\{n_{p}\}$ of spin orbitals $\{|\phi_{p}\rangle\}$ to qubits and nuclear motions to bosonic modes coupled to the qubits, and thus can perform pre-BO dynamics quantum simulations with a complete electronic basis within a given orbital basis set. (d) Existing quantum/classical algorithms rely on BO states $\{|\Phi_{i}\rangle\}$ with potential energy surfaces $\{|E_{i}\rangle\}$ found by solving the electronic time-independent Schrödinger equation using the orbital basis at fixed nuclear positions prior to the simulation. (e) Molecular vibronic dynamics are solved in the BO framework by propagating the nuclear wavefunction (red and blue densities) on a truncated set of coupled BO states.
@@ II.2 Analog quantum simulation
We now introduce our approach to map the Hamiltonian $\hat{H}_{\mathrm{mol}}$ onto an analog quantum device to perform a pre-BO simulation. Such a mapping must satisfy the symmetry requirements of fermions (electrons) and bosons (vibrations), given by their (anti-)commutation relations. Ideally, we would like to map both types of degrees of freedom onto a device with native fermionic [53, 54, 55] and bosonic degrees of freedom. Recently, fermionic quantum processors have been proposed using existing tools such as optical tweezers and neutral fermionic atoms [55]. However, there is currently no quantum architecture which couples fermionic gates to continuous bosonic modes. Therefore, we will build on the wealth of literature for fermion-qubit mappings [56, 57] to propose a digital-analog quantum simulation [58, 59] technique. Our technique takes advantage of existing quantum architectures that have qubit levels and bosonic modes with controllable coupling between them, such as ion traps and circuit quantum electrodynamics (cQED). We will refer to such architectures as “coupled multi-qubit-boson” (cMQB) devices. We note that one could achieve the same type of simulation in a fully analog way if coupled fermionic-bosonic operations are developed in the future.
An illustration of our pre-BO analog simulation method in comparison with the BO framework is summarized in Fig. 1. We start by finding a nuclear position dependent spin orbital basis [Fig. 1(a)] and deriving the expansion coefficients in Eq. (3). We then encode the pre-BO wavefunction [Fig. 1(b)] on a cMQB device, e.g. a trapped-ion device with ion electronic states representing electronic ONVs, and ion motional modes representing the nuclear component of the wavefunction [Fig. 1(c)]. In contrast, existing classical/quantum algorithms in the BO framework require the pre-calculation of BO electronic states [Fig. 1(d)] with the spin orbital basis. The vibronic wavefunction is then propagated on a truncated BO basis [Fig. 1(e)].
@@
The direct implementation of our approach outlined above requires the simultaneous control of all degrees of freedom in the simulation. However, our approach uses a fermion-qubit mapping for the electronic degrees of freedom, which involves a change of basis between non-commuting electronic terms. Therefore, we must use Trotterization [60] of the time-evolution operator. Trotterization approximates the time-evolution operator as a product of operators, each of which corresponds to a term $\hat{H}_{I}$ in Eq. (4). The Trotterized cMQB time-evolution operator is given by $\hat{U}(t)\approx\left(\prod^{N_{\mathrm{op}}}_{I}\hat{U}_{I}(t/N_{\mathrm{t}})\right)^{N_{\mathrm{t}}}\equiv\hat{U}_{\mathrm{cMQB}}(\Delta t)^{N_{\mathrm{t}}}$, where $N_{\mathrm{op}}$ is the number of divided operators and $N_{\mathrm{t}}$ is the number of the Trotter steps [Fig. 2(b)].
@@
where $\hat{\Lambda}^{I}$ is an $N$-qubit Clifford operation which can be composed out of Hadamard, phase, and CNOT gates, to propagate the coupling between bosonic mode(s) and a single qubit $q_{0}^{I}$ generated by a laser-ion interaction to multiple qubits for the desired Pauli strings for $\hat{H}_{I}$. In Fig. 2(b) we show an example time-evolution operator $\hat{U}_{I}(\Delta t)=\exp(-i\theta\Delta t(\hat{b}_{1}+\hat{b}_{1}^{\dagger})\hat{Y}_{2}\hat{Z}_{3}\hat{Y}_{4})$, which can be achieved by a sequence of digital and analog quantum operations. The nuclear-only terms of the Hamiltonian (the final two terms of Eq. (3) are given up to second order by a constant detuning from the bosonic mode frequency. Higher-order nuclear-only terms can be achieved with operations on only the bosonic modes, or qubit-boson operations with an ancilla qubit. The largest coefficients in our Hamiltonian are roughly equal to the orbital energies, thus the maximum Trotter step size $\Delta t$ is inversely proportional to the (active) orbital energy range to achieve the desired Trotter error.
This approach can be considered a digital-analog quantum simulation [58, 59] because the gates involved in generating qubit entanglement take a digital form. Using a coupled fermionic-bosonic operation would eliminate the need for digital gates $\hat{\Lambda}^{I}$ for a fully analog simulation. Even with the fermion-qubit mapping, a more compact form of the time-evolution operator could be achieved using time- and spin-dependent squeezing/displacement operators on the collective ionic motions [61]. For the remainder of this paper we consider the digital-analog approach.
@@
Trapped ions and cQED typically use bosonic modes to achieve entangling gates in digital quantum algorithms. The quantum gates necessary to implement $\hat{\Lambda}^{I}$ are thus readily achieved with these architectures. However, since the entangling gates require a bosonic mode, some bosonic modes (the “bus modes”) must be reserved from those used for simulating nuclear vibrational degrees of freedom. The total number of bosonic modes in the simulator must therefore be at least one greater than the total number of molecular vibrational modes.
@@
As an analog simulation approach, our approach is subject to environmental noise over the course of the simulation, which is a source of error in addition to the Trotter error. The noise effect on an open quantum system can be described by the Lindblad master equation [64]:
@@
where $\hat{\rho}=|\Psi\rangle\langle\Psi|$ represents the density operator of the system, $\hat{H}$ is the Hamiltonian of the system, $D[\hat{L}]\hat{\rho}=\hat{L}\hat{\rho}\hat{L}^{\dagger}-\frac{1}{2}\{\hat{L}^{\dagger}\hat{L},\hat{\rho}\}$ is the Lindblad superoperator, $\hat{L}_{i}$ is the jump operator for the noise source $i$, and $\gamma_{i}$ is the corresponding dissipation rate. The noise effect on a cMQB device translates into the molecular state during a cMQB simulation, where the native rates of the cMQB simulator are scaled to the molecular system depending on the scaling factor $F$ ($\hat{H}_{\mathrm{sim}}=F\hat{H}_{\mathrm{mol}}$), encoding, device parameters, and simulation setup.
@@
where $\Delta t_{\mathrm{mol}}$ is the Trotter step size in the molecular scale, $t_{\mathrm{CNOT}}$ is the experimental time for a two-qubit entangling gate, and $N_{\mathrm{CNOT}}$ is the number of two-qubit entanglement gates per single Trotter step (See Supplemental Material [48]). Although a small Trotter step $\Delta t_{\mathrm{mol}}$ reduces the theoretical Trotter error of the cMQB time-evolution operator, the above equations imply that trade-off exists due to the number of digital qubit entanglement operations.